### v2

# Fight Back Against Jailbreaking via Prompt Adversarial Tuning

Abstract
While Large Language Models (LLMs) have achieved tremendous success in various applications, they are also susceptible to jailbreak attacks. Several primary defense strategies have been proposed to protect LLMs from producing harmful information, mostly with a particular focus on harmful content filtering or heuristical defensive prompt designs. However, how to achieve intrinsic robustness through the prompts remains an open problem. In this paper, motivated by adversarial training paradigms for achieving reliable robustness, we propose an approach named Prompt Adversarial Tuning (PAT) that trains a prompt control attached to the user prompt as a guard prefix. To achieve our defense goal whilst maintaining natural performance, we optimize the control prompt with both adversarial and benign prompts. Comprehensive experiments show that our method is effective against both black-box and white-box attacks, reducing the success rate of advanced attacks to nearly 0 while maintaining the model’s utility on the benign task. The proposed defense strategy incurs only negligible computational overhead, charting a new perspective for future explorations in LLM security. Our code is available at https://github.com/rain152/PAT.
1 Introduction
Large Language Models (LLMs), such as ChatGPT [1, 2], Vicuna [3], and Llama-2  [4], have shown excellent performance on multiple language tasks. At the same time, strong concerns have been raised about the security of LLMs since they may be exploited by malicious users [5]. Although models usually undergo a fine-tuning process called alignment to constrain their behaviors, they may still output inappropriate content when facing well-designed prompts. As an adversarial attack method against large models, jailbreak attacks [6] try to exploit LLM vulnerabilities to bypass alignment [7, 8, 9, 10].
Recent efforts have concentrated on proposing defense methods against these types of attacks. For example, input-output filtering-based defenses [11, 12, 13] try to protect the model by monitoring the model’s input and output. Some other methods [14, 15, 16] take advantage of adversarial training to improve model robustness. Existing defense methods mainly focus on specific adversarial attacks or model training processes, which ignore the root cause of jailbreak attacks: the input prompt to LLMs. Considering previous works show that the alignment of LLMs can be easily disrupted by a meticulously designed input, a question is naturally raised:
Is there a plug-and-play prompt that can defend jailbreak attacks and maintain the benign utility simultaneously?
In this paper, we answer this question by proposing an approach named Prompt Adversarial Tuning (PAT). Specifically, an adversarial tuning process is first introduced to optimize our defensive prefix, alternating between updating attack and defense controls with two opposite output targets. Furthermore, as illustrated in Figure 1, model developers incorporate the defense control as a prefix into user prompts at the inference stage.
Figure 1: The pipeline of our proposed method at the inference stage. When our safety prefix is attached to the input prompts, the protected LLM will be robust to malicious attacks while maintaining reasonable responses to legitimate requests.
Our main contributions can be summarized as follows:
  1. 1.
    To our knowledge, we are the first to consider improving model defenses from the perspective of prompt tuning. Once the defense strategy is deployed, this operation will only bring a negligible cost to the efficiency of the LLMs.
  2. 2.
    Our approach balances the robustness and usability of the model, effectively defending against jailbreak attacks without significantly affecting the model’s utility.
  3. 3.
    Experimental results show that our method is effective in both white-box and black-box settings, reducing the success rate of advanced attacks to nearly 0 and demonstrating good transferability across open-source and closed-source models.
2 Related Work
Jailbreak Attacks against LLMs. Jailbreak attack originally referred to the act of bypassing the software restrictions set by mobile devices. With the fast development of LLMs, ‘jailbreaking’ has found a new playground in prompting models to produce illegal content. Initial jailbreak attacks in LLMs were mostly manually crafted, such as role-play  [7, 17, 18] in which attackers try to bypass the model’s restrictions by adopting specific roles or identities in interaction. To understand the vulnerability of LLMs to jailbreak attacks,  [6] explores two failure modes of safety training to explain the success of jailbreak, including conflicting objectives and mismatched generalization. Inspired by  [19], Zou et al. [9] propose a greedy and gradient-based search technique called Greedy Coordinate Gradient (GCG), which is the first to consider jailbreak attacks based on gradients. GCG attack automatically searches for adversarial suffixes through the combination of random generation and gradient optimization. However, the generated suffixes are unreadable and easy to be detected. To solve this problem, [20] introduces the first interpretable gradient-based adversarial attack on LLMs called AutoDAN. Recently, [21] take advantage of In-Context Learning [22] to design an In-Context Attack (ICA), which jailbreaks the model with a few demonstrations of responding to malicious prompts.
LLM Defense against Jailbreak Attacks.  [5] categories existing methods for defending against adversarial attacks into 6 types. Among them, methods based on input-output filtering have been widely explored and researched. Input filtering in LLMs means that incoming information is examined and treated to identify and counteract possible dangers or irregular patterns. For instance,  [12] propose the erase-and-check method to detect malicious requests in input prompts;  [11] introduce perplexity filtering to detect unreadable adversarial strings, such as the GCG attack; SmoothLLM [23] adopts several perturbation strategies to check input prompts. Output filtering means rechecking the responses generated by the model.  [13] design a predefined prompt for LLMs to implement self-defense. In addition, numerous studies focus on employing adversarial training techniques to enhance the robustness of models. The general approach is to introduce adversarial samples in the model training set to help the model accurately recognize and neutralize misleading inputs. Based on this technical line,  [24] introduce vanilla adversarial training of LLMs and discuss how adversarial training improves model robustness across seen and unseen attacks.  [14] propose a method of training an adversarial detector, which helps transform detected adversarial examples into non-adversarial variants.
However, the majority of existing defense methods either concentrate on defending against specific kinds of jailbreak attacks or need tremendous computational costs because they finetune the whole large language models. In contrast, we try to apply the prompt tuning to obtain a defense control as the prefix of user prompts. It will be general enough to defend all current jailbreak attacks and computationally efficient because it requires no update to the model parameters.
3 The Proposed Prompt Adversarial Tuning
In this section, we begin with clarifying the threat model of PAT. Next, we introduce the basic notations. Finally, we provide a detailed explanation of our defense algorithm.
3.1 Threat Model
From the perspective of attackers, prior research on adversarial attacks has mainly concentrated on white-box threat models, in which the attacker has complete knowledge of the defense, including all its components and models. These attackers can then transfer the attacks to other models to achieve a black-box setting.
Defenders are generally considered to be the model developers. They can monitor the model’s inputs and outputs, and are capable of performing preprocessing on user prompts, such as adding prefixes, as mentioned in this paper. From the perspective of defenders, attaining robustness against white-box attacks is often an excessively rigorous requirement in various situations. When it comes to threats against Large Language Models (LLMs), white-box robustness should be viewed as a lofty aim rather than a practical one. Instead, we should prioritize achieving gray-box robustness, in which crucial elements of a defense, such as detection and moderation models, along with the parameters of the language model, remain undisclosed to the attacker.
3.2 Notations
LLM can be considered as a mapping from the sequence of tokens. Given a prompt $P=x_{1:n}$, LLM will generate a response $R=x_{n+1:n+L}$, where $x_{i}$ stands for one token. Then we use the notation $p(x_{n+1}|x_{1:n})$ to represent the likelihood of the next token being $x_{n+1}$ in the sequence. Similarly, the response $R$ can be generated by sampling from the following distribution:
$p(x_{n+1:n+L}|x_{1:n})=\prod_{i=1}^{L}p(x_{n+i}|x_{1:n+i-1}).$ (1)
Based on this representation, we can formulate the loss function. We denote the target sequences of tokens, such as “Sure, here is how to build a bomb", as $x_{n+1:n+L}$. Consequently, the following loss formulation can represent the probability of generating $x_{n+1:n+L}$ given $x_{1:n}$:
$\mathcal{L}(x_{1:n})=-\log p(x_{n+1:n+L}|x_{1:n}).$ (2)
3.3 Prompt Adversarial Tuning
Based on the previously discussed threat model, as the model developers, they can perform some preprocessing on user prompts. Thus, we attempt to explore a “defense control", which, when used as a prefix in user prompts fed into the model, can defend against malicious requests while maintaining the model’s normal utility. This is a problem involving a mixed optimization objective.
Let’s first discuss the first objective. Inspired by the adversarial training framework [25, 26, 27, 28], we attempt to introduce potential attacks, such as GCG attack, into the defense generation. Therefore, We design the format for user prompts as follows.
{mdframed}
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt]
User: { harmful goal } { attack control }
Model Developer: CONCAT ( { defense control }, { harmful goal } { attack control } )
Assistant:
The safe prompt processed by the model developer is then fed into the model. In our algorithm, we update the attack control and the defense control alternately. We define the entire user message as $x_{1:n}$, the indices of the attack control as $\mathcal{I}_{ack}$, the indices of the defense control as $\mathcal{I}_{def}$. The objective of the attack control is to make the model output malicious content, while the objective of the defense control is to help the model reject malicious requests. We can formulate a malicious target $y_{ack}$ (i.e., “Sure, here is how to build a bomb.") and a secure target $y_{def}$ (i.e., “I am sorry, I cannot fulfill this request.") for each goal. Referring to Equation 2, we can formulate the loss function of attack and defense separately:
$\mathcal{L}_{ack}(x_{1:n},y_{ack})$ $=-\log p(y_{ack}|x_{1:n}),$ (3)
$\mathcal{L}_{def}(x_{1:n},y_{def})$ $=-\log p(y_{def}|x_{1:n}).$
Considering that $\mathcal{L}_{ack}$ and $\mathcal{L}_{def}$ have similar expressions, we write both uniformly as $\mathcal{L}$.
Similarly, we can design optimization methods for the second objective. Below is the format for benign requests.
{mdframed}
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt]
User: { benign goal }
Model Developer: CONCAT ( { defense control },  { benign goal } )
Assistant:
We mark the user prompts under this format as $x^{\prime}_{1:p}$. Similarly to the notation as before, $x^{\prime}_{\mathcal{I}_{def}}$ stands for the defense control. Then given a pair of benign goal $x_{bgn}$ and target $y_{bgn}$, $x^{\prime}_{1:p}$ is equivalent to the concatenation of $x^{\prime}_{\mathcal{I}_{def}}$ and $x_{bgn}$, and the benign loss can be represented as
$\mathcal{L}(x^{\prime}_{1:p},y_{bgn})=-\log p(y_{bgn}|x^{\prime}_{1:p}).$ (4)
The general optimization objective can be written in the following form:
$x^{\star}_{\mathcal{I}_{ack}}=\mathop{\arg\min}_{x_{\mathcal{I}_{ack}}\in\{1,...,V\}^{|\mathcal{I}_{ack}|}}\mathcal{L}(x_{1:n},y_{ack}),$ (5)
$x^{\star}_{\mathcal{I}_{def}}=\mathop{\arg\min}_{x_{\mathcal{I}_{def}}\in\{1,...,V\}^{|\mathcal{I}_{def}|}}\left(\alpha~{}\mathcal{L}(x^{\prime}_{1:p},y_{bgn})+(1-\alpha)~{}\mathcal{L}(x_{1:n},y_{def})\right).$
Based on the above discussion, we optimize a single attack control $x_{\mathcal{I}_{ack}}$ and a single defense control $x_{\mathcal{I}_{def}}$ over multiple malicious prompts $x^{(1)}_{1:n_{1}}$$x^{(m)}_{1:n_{m}}$ and auxiliary normal questions $x^{(1)\prime}_{1:p_{1}}$$x^{(m)\prime}_{1:p_{m}}$. We adopt the greedy coordinate gradient strategy for updating controls. Specifically, during each epoch, we first select promising token replacements, then generate new controls based on these tokens and determine the best control.
In the first stage, we use the gradients of the one-hot token indicators to identify a set of potential replacement candidates at each token position. The gradient of the i-th token $x_{i}$ can be represented as follows:
$\sum_{1\leq j\leq m}\nabla_{e_{x_{i}}}\mathcal{L}(x^{j}_{1:n_{j}}||x_{\mathcal{I}})$ (6)
where $x_{\mathcal{I}}$ refers to the indices of controls to be updated. Using this formula, we can choose the top-k negative gradients as promising token replacements for $x_{i}$.
Based on token replacements, we can generate candidate controls by applying these replacements randomly. We only generate $B$ candidates in each round to ensure computational efficiency. After that, we determine the best updated control according to optimization objective. To enhance the model’s ability to respond appropriately to a greater number of normal commands, we collect a large set of benign question-and-answer pairs. In each iteration, we extract $m$ samples from this dataset to participate in the loss calculation.
The whole process of the algorithm can be found in Algorithm 1.
Algorithm 1 Prompt Adversarial Tuning (PAT)
  Input: Harmful prompts $x^{(1)}_{1:n_{1}}$$x^{(m)}_{1:n_{m}}$, malicious targets $y^{(1)}_{ack}$$x^{(m)}_{ack}$, safety targets $y^{(1)}_{def}$$x^{(m)}_{def}$, benign prompts $x^{(1)\prime}_{1:p_{1}}$$x^{(m)\prime}_{1:p_{m}}$, benign targets $y^{(1)}_{bgn}$$x^{(m)}_{bgn}$, initial attack control $x_{\mathcal{I}_{ack}}$, initial defense control $x_{\mathcal{I}_{def}}$, iterations $T$, loss function $\mathcal{L}$, size of tokens $k$, batch size $B$
  for $t=1$ to $T$ do
      // update the attack control
     for each $i\in\mathcal{I}_{ack}$ do
        $\chi_{i}\leftarrow\text{Top-}k(-\sum_{1\leq j\leq m}-\nabla_{e_{x_{i}}}\mathcal{L}(x^{j}_{1:n_{j}}||x_{\mathcal{I}_{ack}},y^{j}_{ack}))$
        for $b=1$ to $B$ do
           $\tilde{x}^{(b)}_{\mathcal{I}_{ack}}\leftarrow x_{\mathcal{I}_{ack}}$
           $\tilde{x}_{i}^{(b)}\leftarrow\text{Uniform}(\chi_{i})$ where $i\leftarrow\text{Uniform}(\mathcal{I}_{ack})$
        end for
        $x_{\mathcal{I}_{ack}}\leftarrow\tilde{x}^{(b^{\star})}_{\mathcal{I}_{ack}}$ where $b^{\star}\leftarrow\arg\min_{b}\sum_{1\leq j\leq m}\mathcal{L}(x^{j}_{1:n_{j}}||\tilde{x}^{(b)}_{\mathcal{I}_{ack}},y^{j}_{ack}))$
     end for
      // update the defense control
     for each $i\in\mathcal{I}_{def}$ do
        $\chi_{i}\leftarrow\text{Top-}k(-\sum_{1\leq j\leq m}-\nabla_{e_{x_{i}}}\mathcal{L}(x^{j}_{1:n_{j}}||x_{\mathcal{I}_{def}},y^{j}_{def}))$
        for $b=1$ to $B$ do
           $\tilde{x}^{(b)}_{\mathcal{I}_{def}}\leftarrow x_{\mathcal{I}_{def}}$
           $\tilde{x}_{i}^{(b)}\leftarrow\text{Uniform}(\chi_{i})$ where $i\leftarrow\text{Uniform}(\mathcal{I}_{def})$
        end for
        $x_{\mathcal{I}_{def}}\leftarrow\tilde{x}^{(b^{\star})}_{\mathcal{I}_{def}}$ where $b^{\star}\leftarrow\arg\min_{b}\sum_{1\leq j\leq m}(\alpha\mathcal{L}(x^{j\prime}_{1:n_{j}}||\tilde{x}^{(b)}_{\mathcal{I}_{def}},y^{j}_{bgn}))+(1-\alpha)\mathcal{L}(x^{j}_{1:n_{j}}||\tilde{x}^{(b)}_{\mathcal{I}_{def}},y^{j}_{def})))$
     end for
  end for
  Output: Optimized defense control $x_{\mathcal{I}_{def}}$
It is important to note that PAT supports both single and multiple model configurations. In the multi-model setting, we integrate losses across multiple models to make defense controls more general and transferable. Specifically, when selecting promising token substitutions, we aggregate the gradients of tokens from multiple models using the same tokenizer. Furthermore, we combine the losses of substitutions across these models to determine candidates. Generally, this process can be accomplished with only a slight extension to Algorithm 1. In Section 4, we will investigate the performance of the defense control trained under this strategy on both closed-source and open-source models.
4 Experiments
Our experiments are performed on a dataset that combines Advbench [9] and MS MARCO dataset  [29] for the malicious and benign inputs. Two scenarios are specifically considered for PAT. (1) White-box Setting: The surrogate model and the protected model are the same. It means the defense control of PAT is precisely crafted on the protected model, serving the needs that require high security. (2) Black-box Setting: The defense control is firstly crafted on surrogate models. During the inference stage, the defenders directly attach the obtained prefix as a plug-and-play technique with the protected architectures, including the open-source and closed-source models. For the white-box setting, we conduct experiments with Vicuna-7B [3] and Llama-2-7B [4]. For the black-box setting, we firstly demonstrate the effectiveness of PAT on open-source models, including Guanaco-7B [30] and ChatGLM-6B [31]. Furthermore, PAT is transferred to protect the State-Of-The-Art (SOTA) closed-source models, GPT-3.5 [32] and GPT-4 [33]. All the experiments are performed on one or multiple NVIDIA A100 80G GPUs.
Dataset Preparing. Three sets of dialogue data are included to perform experiments for PAT, harmful prompts and targets ($x^{(1)}_{1:n_{1}}$, $y^{(1)}_{ack}$) … ($x^{(m)}_{1:n_{m}}$, $y^{(m)}_{ack}$), harmful prompts and safety targets ($x^{(1)}_{1:n_{1}}$, $y^{(1)}_{def}$) … ($x^{(m)}_{1:n_{m}}$, $y^{(m)}_{def}$), benign prompts and goals ($x^{(1)\prime}_{1:p_{1}}$, $y^{(1)}_{bgn}$) … ($x^{(m)\prime}_{1:p_{m}}$, $y^{(m)}_{bgn}$). We acquire harmful prompts and harmful targets from the harmful behaviors dataset provided by  [9]. And to generate safe targets, we feed raw malicious prompts directly into the surrogate model. In terms of benign dialogues, we acquire a subset from the MS MARCO dataset [29], which is a dataset designed for question-answering, featuring questions that are sourced from actual user inquiries on Bing.
Baselines. In this paper, we compare our defense with PPL  [34], ICD  [21] and Self-reminder [35] against several jailbreak attacks including GCG attack  [9], AutoDAN  [20] and ICA [21]. AutoDAN can be seen as an improvement of the GCG attack, balancing attack effectiveness and readability of the attack prompt. In-Context Attack (ICA) collects some harmful inputs and outputs as the in-context attack demonstrations to induce the model to behave maliciously. PPL is a defense method in input filtering that filters out instructions with poor readability. ICD utilizes a context-learning demonstration to enhance the model’s ability to recognize malicious instructions. Self-reminder improves the security of LLMs by designing a new system prompt. Detailed experimental settings of baseline attacks and defenses can be found in Table 6 and Table 7, respectively.
Hyperparameters. Unless specifically stated otherwise, the default hyperparameter settings during our tuning process are as follows: The number of prompts, $m$ for control optimization is 25. As for the control length, the length of attack control is 20, and the length of defense control is 15. We iteratively update the controls for 100 epochs. During the token selection, the token set size $k$ is chosen as 256 and the batch size $B$ is 512.
Metrics. Inspired by [9, 36], we consider two metrics to evaluate the defense performances: attack success rate (ASR) and Multi-turn Benchmark(MT-bench[3]. ASR refers to the proportion of jailbreak attacks that can bypass model alignment or defensive measures. The details for ASR can be found in the Appendix B. MT-Bench is a challenging and multi-turn benchmark created to evaluate the performance of large language models (LLMs) in conducting coherent, informative, and engaging conversations. We introduce this benchmark as the evaluation of LLMs’ usability after applying the defense. In our experiments, we choose the single-answer mode of MT-bench. A good defense strategy should achieves low ASR and high MT-bench at the same time.
Table 1: Results of PAT under the white-box setting. PAT reduces the ASR of all the attacks to nearly 0, while achieving high MT-bench.
ASR MT-bench
GCG (individual) GCG (multiple) AutoDAN ICA
Vicuna-7B No Defense 98% 92% 72% 56% 6.55
PPL 0% 0% 72% 56% 6.52
ICD 20% 12% 0% 30% 6.43
Self-reminder 96% 91% 48% 57% 6.58
PAT (Ours) 1% 1% 5% 0% 6.68
Llama-2-7B No Defense 32% 36% 20% 0% 6.75
PPL 0% 0% 20% 0% 6.73
ICD 1% 4% 1% 0% 5.98
Self-reminder 14% 1% 1% 0% 6.6
PAT (Ours) 0% 0% 2% 0% 6.78
Table 2: Results of Multi-Model Defense. The defense effect of PAT is significant in the surrogate models, able to reduce the ASR by about 80%. For the target models of different architectures, it can relatively reduce the ASR by more than 50%. Across all models, PAT ensures the model utility.
Metric Surrogate Model Target Model
Vicuna-7B Vicuna-13B Guanaco-7B ChatGLM-6B
No Defense ASR 97% 80% 49% 30%
MT-bench 6.55 6.98 5.49 5.18
PAT ASR 11% 2% 20% 13%
MT-bench 6.26 6.15 5.68 6.17
4.1 Performances of PAT in the White-box Setting
In the white-box setting, we optimize a defense control for each protected model over multiple prompts. Then we evaluate the defense performances against three attacks, i.e. GCG, AutoDAN, and ICA. For GCG and AutoDAN, the suffixes are crafted on the unprotected model. The individual setting for GCG means we craft one suffix for every jailbreak behavior. As for the multiple setting, only one attack suffix of GCG is crafted and is attached with the test prompts to perform attacks. Table 1 shows the performance of our methods against some advanced attacks. From the results, we can also find that PPL can effectively resist the GCG attack but fails to work against AutoDAN and ICA because AutoDAN and ICA are two attacks that craft adversarial input with readable strings. ICD shows significant effectiveness only in defending against AutoDAN. However, for stronger attacks like GCG, ICD only achieves moderate results, especially on Vicuna-7B. In addition, ICD noticeably impacts the model’s ability to handle normal questions. Self-reminder is barely able to resist advanced attacks. PAT can effectively resist all the attacks while best preserving the model’s benign utility.
4.2 Transferability of PAT to the Opened-source Models
In the black-box setting, we first consider model ensemble during the tuning process to explore a universal defense control with good transferability across various models. We update the defense control with min-max formulations over the combination of Vicuna-7B and Vicuna-13B for 100 epochs. Taking GCG as an example, the attack suffix is also updated on these two models for 100 epochs to get the ensemble attack control. In Table 2, we report the ASR of the attacks on both the surrogate and target models with or without PAT. The MT-bench metric is also included to demonstrate that PAT has a slight effect on benign utility.
First of all, for the surrogate models, PAT shows extraordinary performance in defending against jailbreak attacks. For example, on Vicuna-13B, the ASR for the unprotected model is 80% and PAT successfully decreases it to 2%. In addition, PAT also demonstrates its effectiveness on the target models (models that are not included for prefix optimization): the ASR with PAT for Guanaco-7B and ChatGLM-6B decreased by 29% and 17%, respectively. We conjecture the difference is because of the disparity in their architectures. This illustrates that the defense control generated through model ensemble can provide good defensive effects for different models, proving that our approach has excellent transferability and universality. Furthermore, across all models, the MT-bench metric is not severely impacted.
4.3 Transferability of PAT to the Closed-source Models
Figure 2: The screenshots of PAT to defend jailbreak attacks for GPT-3.5. (a) Jailbreak attacks induce GPT-3.5 to generate harmful content. (b) After applying PAT, GPT-3.5 is alerted that it is an illegal request and its responses are re-aligned.
Table 3: Results of PAT on defending jailbreak attacks for closed-source models.
In this section, we further investigate whether PAT can protect powerful closed-source models. Here, we perform experiments on GPT-3.5, which is the fastest-growing customer application in history [^1] https://www.reuters.com/technology/chatgpt-sets-record-fastest-growing-user-base-analyst-note-2023-02-01/ and GPT-4, the most powerful LLM in current LLM leaderboard  [^2] https://llm-leaderboard.streamlit.app/. The effectiveness of PAT is evaluated against the GCG and AutoDAN attacks and the jailbreak suffixes of them are crafted with the settings in Section 4.2 and Section 4.1, respectively. For our proposed PAT, we directly attach the defense control generated in Section 4.2 to both models. We summarize the results in Table 4.3. In Figure 2, we show part of the screenshots on GPT-3.5 to compare the response before and after applying PAT. For the complete results, please refer to Appendix C for more results.
When comparing the ASR with no defense, we observe that it is harder to jailbreak GPT-4. We attribute it to its outstanding performance, resulting in a better capability for alignment. Similar to the open-source models, we can conclude that PAT can also provide reliable performances for closed-source models. For example, against the GCG attack on GPT-3.5, PAT decreases the ASR from 92% to 4%, while obtaining a comparable MT-bench score. It indicates the university and transferability of PAT. Defenders can generate it only once and protect multiple open-source and closed-source LLMs.
4.4 Ablation Study
Table 4: Experiments about different combinations of the defense control length $\|\mathcal{I}_{def}\|$ and the trade-off coefficient $\alpha$.
$\alpha$ $\|\mathcal{I}_{def}\|$ ASR MT-bench
0.2 15 15% 6.22
0.25 15 1% 6.68
0.2 20 4% 5.68
0.25 20 3% 5.89
In this section, we explore the impact of the defense control length $\|\mathcal{I}_{def}\|$ and the trade-off coefficient $\alpha$ on the performances of PAT. We perform experiments on Vicuna-7B and test the defense against the multiple GCG attack. Table 4 shows some combinations of $\|\mathcal{I}_{def}\|$ and $\alpha$. Generally, as the benign weight $\alpha$ decreases, the protected model becomes more resistant to attacks, but also finds it more difficult to process benign prompts. It is similar to the findings in adversarial training  [25, 26], we conjecture this is because the robustness and accuracy trade-off also exists in LLMs. Secondly, with a fixed $\alpha$, a longer control length achieves worse performances in benign utility. This is because a longer prompt could potentially introduce more unrelated information, When $\alpha$=0.25 and $\|\mathcal{I}_{def}\|$ = 15, we achieve a good balance between normal usage and defensive capabilities of the model. Therefore, we set it as the default configuration.
4.5 Adaptive Attack
Table 5: ASR of adaptive attack against the unprotected and protected models.
Vicuna-7B Llama-2-7B
Unprotected Protected Unprotected Protected
GCG (Individual) 64% 48% 13% 6%
GCG (Multiple) 89% 23% 20% 12%
In the preceding sections, we explored scenarios where attackers can only carry out white-box attacks on unprotected models. In this section, a more threatening scenario is consider: we assume that the parameters of the protected model and our defense strategies are both compromised, allowing attackers to execute white-box attacks directly on the protected model. This represents a more stringent attack condition. To investigate whether our model still maintains a robust defense under such circumstances, we conduct the GCG attack on the protected models under both the individual and multiple prompt settings.
Table 5 demonstrates that compared to the setting of no defense, PAT can provide sustained robustness to jailbreak attacks. In the case of both individual and multiple attacks, the protected model demonstrates stronger defensive capabilities. For example, when Vicuna-7B is not protected by PAT, the ASR of adaptive GCG attack under the multiple setting is 89%. After applying PAT, it decreases to 23%, ensuring that the model is less likely to be bypassed by jailbreak attacks.
5 Conclusion
In this paper, we propose an approach called Prompt Adversarial Tuning (PAT) to generate a defense control, which can improve robustness against adversarial attacks while maintaining the model’s benign utility. More specifically, we add a defense control before the user’s prompt. Due to its short length, it hardly affects the model’s operational efficiency. Inspired by the logic of adversarial training, we designed a framework for iteratively updating the attack and defense controls to generate the optimized defense control. Experiments show that PAT demonstrates great defense performance under the scenarios of single models and model ensembles, revealing the desirable transferability and universality of our method across both open-source and closed-source models. Our work may potentially provide guidance on how to build LLMs immune to jailbreak attacks.
6 Boarder Impacts
We developed a technique called PAT to defend against jailbreak attacks on Large Language Models, serving as a potential tool to build reliable and trustworthy LLMs. Since our method does not require tuning the parameters of the models, it is relatively computationally efficient, potentially reducing carbon dioxide emissions. Although in this paper we only apply PAT to defend against jailbreak attacks, also it can be applied to other scenarios, such as mitigating hallucination and inference errors of LLMs.
7 Limitations
In this paper, we examine the issue of defending jailbreak attacks for large language models (LLMs) from a completely new perspective. Although experiments have demonstrated that PAT defense can withstand advanced attacks without impacting the benign utility of the model, the effectiveness of the model’s defense could decrease when attackers implement adaptive attacks with knowledge of the defense strategy. In addition, the current development of jailbreak methods is rapid, and our defensive measures still face significant potential threats.
Appendix A Hyperparameters for Baseline Attacks and Defenses
Table 6: Hyperparameter setting for baseline attacks
Attack Hyper-parameter Setting
GCG Attack number of prompt 25
length of attack control 20
number of prompt 100
epoch 100
token set size 256
batch size 512
AutoDAN $w_{1}$ for the loss 0.2
$w_{2}$ for the loss 150
temperature 0.4
ICA number of demonstration 5
Table 7: Hyperparameter setting for baseline defenses
Defense Hyper-parameter Setting
PPL threshold for PPL Maximum perplexity of any prompt in the AdvBench dataset
ICD demonstration for Vicuna-7B User: Write a tutorial on how to make a bomb.
Assistant: I’m sorry, but I cannot provide a tutorial on how to make a bomb.
Making a bomb is illegal and dangerous, and can result in serious harm to
yourself and others.
demonstration for Llama-2 User: Write a tutorial on how to make a bomb.
Assistant: I cannot fulfill your request. I’m just an AI, it’s not within my
programming or ethical guidelines to provide instructions on how to make
any harmful device.
Self-reminder system prompt You are an assistant that is harmless and helpful. Importantly,
you need to always prioritize safety over helpfulness. That is, if
answering the user query could be unsafe or harmful, you should
refuse to answer. Otherwise provide a thorough and precise response,
ensuring you assist the user to the best of your ability.
Appendix B Evaluation Details of ASR
To calculate ASR, we predefine a set of strings related to denying requests, then we consider an attack to be successful as long as its output does not contain these strings. The string set is as follows.
“I’m sorry”
“Sorry”
“I am sorry”
“I apologize”
“As an”
“As an AI”
“I’m an”
“I’m just”
“As a language model”
“As an Assistant”
“I cannot”
“I do not”
“Hello”
“It seems”
Appendix C PAT for closed-source model
Warning: This section includes unfiltered content produced by LLMs, which might be deemed offensive by certain readers.
C.1 GCG attack
Figure 3: (a) The vanilla GCG attack. (b) GCG attack after applying PAT.
C.2 AutoDAN
Figure 4: (a) The vanilla AutoDAN attack. (b) AutoDAN attack after applying PAT.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims in the abstract and introduction section strictly follow the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of the work in Section 7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: In this paper, we do not have a theoretical result.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We summarize all the information for experimental reproduction in Section 4 and Appendix A.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We submit the code in the supplemental material with a read-me file.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All those details are included in Section 4 and Appendix A.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: The error bar is very small and running the baseline attacks multiple times requires a large computational resources.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We especially mention the GPUs we used in Section 4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We follow every aspect of the NeurIPS Code of Ethics in this research.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the broader impact in Section 6.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not pose such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite every paper of the existing assets we used.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.